As studied, the average number of hours spent by senior high school students for their online classes a week is 25 hours with a standard deviation of 4 hours. Assuming that study is true and the data is normally distributed.


Find the probability that a randomly selected senior high student spends more than 26 hours but less than 29 hours.

1
Expert's answer
2022-02-27T17:22:19-0500

we use z scores

z(26) = (26 - 25)/ 4 = 0.25

z(29) = (29-25) /4 = 1

we define p(0.25<z<1), using the normal tables we obtain

(0.84134 - 0.59871) = 0.24263 which is the required solution.